Calculate the molarity of each of the following:
1. 66.5 g LiCl in 250.0 mL of solution
2. 30.5 g NiSO4 in 200.0 mL of solution
3. 51.0 g KCN in 500.0 mL of solution
1
Expert's answer
2019-11-18T06:28:50-0500

Molarity,"M=n\/v" . Where n is the number of moles and v is the volume in litres.

Number of moles, "n =mass\/RFM."

1.      66.5g LiCl= (66.5/42.4) =1.568 moles

Hence M=1.568/0.25 =6.272 mol/lit

2.      30.5g NiSO4= (30.5/154.8) = 0.197 moles

Hence M =0.1970/0.2 = 0.985 mol/lit

3.      51.0g KCN = (51/65.1) =0.783 moles

Hence M=0.783/0.5 = 1.566 mol/lit